OpenStudy (anonymous):

write the equation of the line that passes through (2,5) with a slope of 3 in point slope form

OpenStudy (ivettef365):

you need to use the formula y = mx + b

OpenStudy (anonymous):

ok and what do i do after i have the formula

OpenStudy (anonymous):

i really dont understand this norr know how to use the formula so plz help

OpenStudy (ivettef365):

this is how you read the formula y = mx + b y = a point x = a point m = slope b = y- int

OpenStudy (ivettef365):

so you have slope = 3 = m, points (2,5) which is (x,y), now you need to find the value of b

OpenStudy (ivettef365):

can you do that

OpenStudy (anonymous):

no not really

OpenStudy (anonymous):

wouold the 5 replace the y ?

OpenStudy (ivettef365):

well just plug in the points and the m (slope) as follows: y = mx + b 5 = 3(2) + b can you solve for b

OpenStudy (anonymous):

would it be -1

OpenStudy (ivettef365):

correct, now when they ask you the equation of the line, what the question is asking you is for the formula of y = mx + b, then only thing you need to find is the m and the b

OpenStudy (ivettef365):

so since you have m = 3 and b = -1 then the equation of the line would be y = 3x -1

OpenStudy (ivettef365):

the reason why you don't have values for (x,y), because a line have many points, so the equation of the line (not the point) is written as y = 3x -1, when you have a slope of 3 and passes through those points

OpenStudy (anonymous):

o ok .. thank you so much i would have never know what to do if it wasnt for you